# On the semi-additivity of the $1/2$-symmetric caloric capacity in the plane

### v3

@@ 6.2. General comparability for parallelepipeds in $\mathbb{R}^{n+1}$
Observe that in the proofs of Lemma 6.4 and Corollary 6.5 the constants appearing in $\widetilde{\gamma}(R)\approx\widetilde{\gamma}_{+}(R)$ and $\widetilde{\gamma}(E)\approx\widetilde{\gamma}_{+}(E)$ do not depend on the minimal side length $\ell_{\text{min}}$ of the parallelepiped $R$ in the first case, nor on $\delta>0$ in the second; both being strictly positive parameters. Such quantities, however, do determine the number of steps needed to carry out the iterative argument of the proofs. But this is not an issue, since the constant appearing in assumption $\text{{{A}}}_{\boldsymbol{3}}$ is precisely taken to be $C_{1}$, the same constant of property $\text{{{P}}}_{\boldsymbol{3}}$ in Theorem 4.1. This enables to carry out the estimates of (6.2) and avoid any possible dependence on $\ell_{\text{min}}$ or $\delta$ when tracing back each step of the iteration, since $C_{1}$ cancels itself out with its own inverse. Again, let us remark that this type of argument has been inspired by that of [Vo, §6].
Let us apply Corollary 6.5 to the usual family of corner-like Cantor sets of $\mathbb{R}^{n+1}$. Let $\lambda=(\lambda_{j})_{j}$ be a sequence of real numbers satisfying $20<\lambda_{j}<1/$ 2. Define its associated Cantor set $E\subset\mathbb{R}^{n+1}$ by the following algorithm. Set $Q^{0}:=[0,1]^{n+1}$ the unit cube of $\mathbb{R}^{n+1}$ and consider $2^{n+1}$ disjoint cubes inside $Q^{0}$ of side length $\ell_{1}:=\lambda_{1}$, with sides parallel to the coordinate axes and such that each cube contains a vertex of $Q^{0}$. Continue this same process now for each of the $2^{n+1}$ cubes from the previous step, but now using a contraction factor $\lambda_{2}$. That is, we end up with $2^{2(n+1)}$ cubes with side length $\ell_{2}:=\lambda_{1}\lambda_{2}$. Proceeding inductively we have that at the $k$-th step of the iteration we encounter $2^{k(n+1)}$ cubes, that we denote $Q_{j}^{k}$ for $1\leq j\leq 2^{k(n+1)}$, with side length $\ell_{k}:=\prod_{j=1}^{k}\lambda_{j}$. We will refer to them as cubes of the $k$-th generation. We define
$E_{k}=E(\lambda_{1},\ldots,\lambda_{k}):=\bigcup_{j=1}^{2^{k(n+1)}}Q_{j}^{k},$
and from the latter we obtain the Cantor set associated with $\lambda$,
$E=E(\lambda):=\bigcap_{k=1}^{\infty}E_{k}.$
If we chose $\lambda_{j}=2^{-(n+1)/n}$ for every $j$ we would recover the particular Cantor set presented in [MPr, §5]. Let us introduce the following density for each $k\geq 1$,
$\theta_{k}:=\frac{2^{-k(n+1)}}{\ell_{k}^{n}},$
where $Q^{k}$ is any cube of the $k$-th generation. We also set $\theta_{0}:=1$. Combining Theorem 3.5, the results of [He, §4] and Corollary 6.5 we deduce:
Corollary 6.6.
Let $(\lambda_{j})_{j}$ be a sequence of real numbers satisfying $20<\lambda_{j}\leq\tau_{0}<1/$ 2, for every $j$. Then,
$\widetilde{\gamma}(E_{k})\approx\widetilde{\gamma}_{+}(E_{k})\approx\bigg{(}\sum_{j=0}^{k}\theta_{j}\bigg{)}^{-1},$
where the implicit constants only depend on $n$ and $\tau_{0}$. Moreover,
$\widetilde{\gamma}(E)\approx\widetilde{\gamma}_{+}(E)\approx\bigg{(}\sum_{j=0}^{\infty}\theta_{j}\bigg{)}^{-1}.$
Proof.
The result for $E_{k}$ is a direct application of [He, Theorems 4.1 & 4.2], Theorem 3.5 and Corollary 6.5. To obtain the result for $E$ notice that for each generation $k$ we have,
$\widetilde{\gamma}(E)\leq\widetilde{\gamma}(E_{k})\approx\widetilde{\gamma}_{+}(E_{k})\leq\gamma_{+}(E_{k})\approx\bigg{(}\sum_{j=0}^{k}\theta_{j}\bigg{)}^{-1}.$
Therefore, using the outer regularity of $\gamma_{+}$ [He, Proposition 3.1] we get
$\widetilde{\gamma}(E)\lesssim\gamma_{+}(E)\approx\bigg{(}\sum_{j=0}^{\infty}\theta_{j}\bigg{)}^{-1},$
and by Theorem 3.5 we are done. ∎
6.3. General comparability for compact sets in $\mathbb{R}^{2}$
This subsection deals with the proof of the following result:
@@ 6.3. General comparability for compact sets in $\mathbb{R}^{2}$
Let $E\subset\mathbb{R}^{2}$ be a compact set. Then,
@@
Let $E\subset\mathbb{R}^{2}$ be a compact set, that we assume without loss of generality that satisfies assumptions $\text{{{A}}}_{\boldsymbol{1}}$ and $\text{{{A}}}_{\boldsymbol{2}}$. So in particular $E$ is contained in the unit ball and consists of a finite union of dyadic cubes belonging to a dyadic grid in $\mathbb{R}^{2}$ ((((with sides parallel to the coordinate axes))\,)), all of the same size and with disjoint interiors. We denote by
[…]
Again, we repeat the iterative scheme of Lemma 6.4: in general, at step $S\geq 1$ of the iteration we would obtain for a set of the form
$E_{i_{1}i_{2}\cdots i_{S-1}}=E,\qquad\text{if}\quad S=1,$
$E_{i_{1}i_{2}\cdots i_{S-1}}=2\pazocal{Q}_{i_{1}i_{2}\cdots i_{S-1}}\cap E_{i_{1}i_{2}\cdots i_{S-2}},\qquad\text{if}\quad S>1,$
a family of cubes $\big{\{}\pazocal{Q}_{i_{1}i_{2}\cdots i_{S-1}1},\ldots,\pazocal{Q}_{i_{1}i_{2}\cdots i_{S-1}N_{i_{1}i_{2}\cdots i_{S-1}}}\big{\}}$ with diameters comparable to $5^{-S}\text{diam}(E)$. Now we would distinguish the cases
  1. A)
    If $E_{i_{1}i_{2}\cdots i_{S}}:=2\pazocal{Q}_{i_{1}i_{2}\cdots i_{S}}\cap E_{i_{1}i_{2}\cdots i_{S-1}}$ has diameter smaller than $\delta/4$ for every $i_{S}=1,\ldots,N_{i_{1}i_{2}\cdots i_{S-1}}$.
  2. B)
    Or if there exists an index $i_{S}$ so that $\text{diam}(E_{i_{1}i_{2}\cdots i_{S}})\geq\delta/4$.
As well as
  • 1)
  • 2)
If 1) occurs we are done, in the sense that we can deduce the estimate
$\widetilde{\gamma}(E_{i_{1}i_{2}\cdots i_{S-1}})\lesssim\widetilde{\gamma}_{+}(E_{i_{1}i_{2}\cdots i_{S-1}}).$
If 2) occurs, we move on to the next step of the iteration up to the point where $S$ is large enough (depending on $\delta$) so that option B) is no longer possible. That is, the iteration stops once the only two possible scenarios are: 1) or 2)A). In this setting, we are left to study the case where 2) and A) occur simultaneously. To deal with this case, we write the precise definition of $E_{i_{1}i_{2}\cdots i_{S}}$, which is
$E_{i_{1}i_{2}\cdots i_{S}}:=\big{(}2\pazocal{Q}_{i_{1}i_{2}\cdots i_{S}}\cap 2\pazocal{Q}_{i_{1}i_{2}\cdots i_{S-1}}\cap\cdots\cap 2\pazocal{Q}_{i_{1}i_{2}}\cap 2\pazocal{Q}_{i_{1}}\big{)}\cap E=:2R_{i_{1}i_{2}\cdots i_{S}}\cap E,$
where $R_{i_{1}i_{2}\cdots i_{S}}$ is a rectangle with diameter comparable to $5^{-S}\text{diam}(E)\ll\delta$. In fact, applying properties $\text{{{P}}}_{\boldsymbol{1}}$ and $\text{{{P}}}_{\boldsymbol{4}}$ in Theorem 4.1 at each step, one easily deduces the inclusions
(6.4) $\frac{11}{16}\pazocal{Q}_{i_{1}i_{2}\cdots i_{S}}\subset 2R_{i_{1}i_{2}\cdots i_{S}}\subset 2\pazocal{Q}_{i_{1}i_{2}\cdots i_{S}}.$
The latter is trivial for $S=1$. If $S>1$, the argument is similar to one that is presented in the proof of Lemma 6.4. For example, if $S=2$, $\text{{{P}}}_{\boldsymbol{1}}$ yields, for each $i_{2}=1,\ldots,N_{i_{1}}$,
$\frac{5}{8}\pazocal{Q}_{i_{1}i_{2}}\cap(2\pazocal{Q}_{i_{1}}\cap E)\neq\varnothing,\qquad\text{so in particular}\qquad\frac{5}{8}\pazocal{Q}_{i_{1}i_{2}}\cap 2Q_{i_{1}}\neq\varnothing.$
On the other hand, $\text{{{P}}}_{\boldsymbol{4}}$ yields $\text{diam}(\pazocal{Q}_{i_{2}i_{2}})\leq\frac{1}{10}\text{diam}(2\pazocal{Q}_{i_{1}}\cap E)$, which implies
$\ell(\pazocal{Q}_{i_{1}i_{2}})\leq\frac{1}{5}\ell(\pazocal{Q}_{i_{1}}),\qquad\text{for each}\quad i_{2}=1,\ldots,N_{i_{1}}.$
All in all, one easily gets $\frac{11}{16}\pazocal{Q}_{i_{1}i_{2}}\subset 2\pazocal{Q}_{i_{1}i_{2}}\cap 2\pazocal{Q}_{i_{1}}=:R_{i_{1}i_{2}}$, which are the desired inclusions. Such scheme can be repeated at each step to obtain, in general, relation (6.4). The latter inclusions imply that the rectangle $R_{i_{1}i_{2}\cdots i_{S}}$ behaves as a square, in the sense that its side lengths are comparable and thus, by the second statement in Corollary 6.3,
$\widetilde{\gamma}(R_{i_{1}i_{2}\cdots i_{S}})\leq 2\,\gamma(\pazocal{Q}_{i_{1}i_{2}\cdots i_{S}})\approx\gamma_{+}(\pazocal{Q}_{i_{1}i_{2}\cdots i_{S}})\leq\frac{16}{11}\,\gamma_{+}(R_{i_{1}i_{2}\cdots i_{S}})\approx\widetilde{\gamma}_{+}(R_{i_{1}i_{2}\cdots i_{S}}).$
That is, $\widetilde{\gamma}(R_{i_{1}i_{2}\cdots i_{S}})\approx\widetilde{\gamma}_{+}(R_{i_{1}i_{2}\cdots i_{S}})$. In addition, since A) occurs, by the definition of $\delta$ we get that $2R_{i_{1}i_{2}\cdots i_{S}}$ can only intersect one connected component of $E$. Moreover, if $2R_{i_{1}i_{2}\cdots i_{S}}$ intersects one of the dyadic squares that conform $E$, it can only intersect, in addition, those squares which are adjacent to it (a maximum of four). Therefore, there is only a finite number of possible compact sets that $E_{i_{1}i_{2}\cdots i_{S}}:=2R_{i_{1}i_{2}\cdots i_{S}}\cap E$ can be. We represent them in Figure 1. There, $2R_{i_{1}i_{2}\cdots i_{S}}$ is assumed, without loss of generality by (6.4), to be a (red) square; and some of the dyadic squares conforming $E$ are depicted in gray.
Figure 1. In black, the eight possible intersections that can be obtained from $E_{i_{1}i_{2}\cdots i_{S}}:=2R_{i_{1}i_{2}\cdots i_{S}}\cap E$, for $S$ large enough so that $5^{-S}\text{diam}(E)\ll\delta$. In red the rectangle $2R_{i_{1}i_{2}\cdots i_{S}}$ (assumed to be a square by relation (6.4)), and in gray some of the dyadic squares conforming $E$. Let us remark that in $d)$ we also contemplate the case where $2R_{i_{1}i_{2}\cdots i_{S}}\cap E=2R_{i_{1}i_{2}\cdots i_{S}}$.
It is clear that in cases a), b), c) and d), a direct application of Lemma 6.4 yields
$\widetilde{\gamma}(E_{i_{1}i_{2}\cdots i_{S}})\lesssim\widetilde{\gamma}_{+}(E_{i_{1}i_{2}\cdots i_{S}}).$
Regarding f), g) and h), we write $2\ell^{H}_{i_{1}i_{2}\cdots i_{S}}$ the horizontal side length of $2R_{i_{1}i_{2}\cdots i_{S}}$, and $I^{H}_{i_{1}i_{2}\cdots i_{S}}$ the horizontal side of $E_{i_{1}i_{2}\cdots i_{S}}$ with such length. Then,
$\widetilde{\gamma}(E_{i_{1}i_{2}\cdots i_{S}})\leq\widetilde{\gamma}(2R_{i_{1}i_{2}\cdots i_{S}})\approx\widetilde{\gamma}_{+}\big{(}I^{H}_{i_{1}i_{2}\cdots i_{S}}\big{)}\leq\widetilde{\gamma}_{+}(E_{i_{1}i_{2}\cdots i_{S}}),$
where for the second inequality we have used the second statement in Corollary 6.3 and relation (6.4), and the third follows simply by the monotonicity of the capacity. Hence we also obtain the desired estimate. Finally, case e) can be dealt with in a similar way, just noticing that at least one of the diagonally opposed rectangles $R_{i_{1}i_{2}\cdots i_{S}}^{\uparrow}$ and $R_{i_{1}i_{2}\cdots i_{S}}^{\uparrow}$, obtained in $E_{i_{1}i_{2}\cdots i_{S}}$, presents horizontal side length half of $2\ell^{H}_{i_{1}i_{2}\cdots i_{S}}$. Assume it is $R_{i_{1}i_{2}\cdots i_{S}}^{\uparrow}$. Then,
$\widetilde{\gamma}(E_{i_{1}i_{2}\cdots i_{S}})\leq\widetilde{\gamma}(2R_{i_{1}i_{2}\cdots i_{S}})\approx\widetilde{\gamma}_{+}\big{(}I^{H}_{i_{1}i_{2}\cdots i_{S}}\big{)}\leq 2\,\widetilde{\gamma}_{+}\big{(}R_{i_{1}i_{2}\cdots i_{S}}^{\uparrow}\big{)}\leq 2\,\widetilde{\gamma}(E_{i_{1}i_{2}\cdots i_{S}}).$
Therefore, in all possible scenarios we get
$\widetilde{\gamma}(E_{i_{1}i_{2}\cdots i_{S}})\lesssim\widetilde{\gamma}_{+}(E_{i_{1}i_{2}\cdots i_{S}}),$
so applying the same type of estimates of (6.2) (since we are assuming that $\text{{{A}}}_{\boldsymbol{3}}$ is not satisfied) we deduce
$\widetilde{\gamma}(E_{i_{1}i_{2}\cdots i_{S-1}})\lesssim\widetilde{\gamma}_{+}(E_{i_{1}i_{2}\cdots i_{S-1}}),$
the necessary estimate to trace back the iterations and obtain the desired result. ∎
@@
The capacity $\widetilde{\gamma}$ is semi-additive in $\mathbb{R}^{2}$. That is, there is an absolute constant $C>0$ so that for any $E_{1},E_{2},\ldots$ disjoint compact sets of $\mathbb{R}^{2}$,
@@
In this section we compute the $\widetilde{\gamma}$ capacity of a closed rectangle $R\subset\mathbb{R}^{2}$ with sides parallel to the coordinate axes and respective side lengths $\ell_{x}>0,\,\ell_{t}>0$. More precisely, we obtain the following result:
@@
We introduce the parameter
$r:=\frac{\ell_{x}}{\ell_{t}},$
that is nothing but the spatial side length of $R_{0}$, as well as the measure
$\mu:=\pazocal{L}^{2}|_{R_{0}}.$
@@
it follows that for any fixed $t$,
$P\ast\mu(x,t)=P\ast\mu(r-x,t),$
or in other words, the one variable function $x\mapsto P\ast\mu(x,t)$ is symmetric with respect to the point $x=r/2$, for any $t\in\mathbb{R}$. Moreover, it is clear that it is nonnegative, tends to 0 as $x\to\pm\infty$ and in fact, for $t>0$, it attains its maximum precisely at $x=r/2$. This last property can be argued as follows: begin by fixing $t\in(0,1]$ and noticing that
@@
[…]
@@
Using the particular value of $M(r):=P\ast\mu(r/2,1)$ we are able to obtain a lower bound for $\gamma_{+}$:
@@
Therefore, regarding the original rectangle $R$ and applying Theorems 3.5 and 6.7, Theorem 7.1 follows. ∎

### v4

@@ 6.2. General comparability in $\mathbb{R}^{n+1}$
Observe that constants in Lemmas 6.4 and 6.5 do not depend on $\ell_{\text{min}}$ in the first case, nor on $\delta>0$ in the second. Such positive quantities, however, do determine the number of steps needed to carry out the iterative argument of the proofs. But this is not an issue, since the constant in $\text{{{A}}}_{\boldsymbol{3}}$ is taken to be $C_{1}$ and it cancels with its own inverse from property $\text{{{P}}}_{\boldsymbol{3}}$ in Theorem 4.1, eliminating possible dependence on $\ell_{\text{min}}$ or $\delta$ when tracing back each step of the iteration.
Remark 6.3.
In Lemma 6.5 we have claimed that $\widetilde{\gamma}_{+}(E)\lesssim\widetilde{\gamma}_{+}(\mathring{E})$, for $E=Q_{1}\cup\cdots\cup Q_{N}$ a union of $N$ cubes with disjoint interiors. We will give the scheme to prove this by induction. The case $N=1$ is clear: first notice that $\widetilde{\gamma}(Q_{1})\lesssim\pazocal{H}^{n}_{\infty}\leq\ell(Q_{1})^{n}$. Now, pick any $0<\lambda<1$ and set $Q_{1}^{\lambda}\subset\mathring{Q_{1}}$ the closed cube concentric with $Q_{1}$ and side length $\lambda\ell(Q_{1})$. Consider $\mu^{\lambda}:=\ell(Q_{1}^{\lambda})^{-1}\pazocal{L}^{n+1}|_{Q_{1}^{\lambda}}$ and observe that $P\ast\mu^{\lambda}\lesssim 1$ and $P^{\ast}\ast\mu^{\lambda}\lesssim 1$. Then, $\widetilde{\gamma}_{+}(Q^{\lambda}_{1})\geq\mu^{\lambda}(Q_{1}^{\lambda})=\ell(Q_{1}^{\lambda})^{n}$, and making $\lambda\to 1$ we get $\widetilde{\gamma}_{+}(\mathring{Q}_{1})\geq\ell(Q_{1})^{n}$, so the case $N=1$ is proved. Now, we proceed by induction assuming the case $N$ true and proving the case $N+1$. To do so, we simply use the semi-additivity of $\widetilde{\gamma}_{+}$ (see [He, Proposition 3.1, Theorem 3.11]) followed by the induction hypothesis:
$\widetilde{\gamma}_{+}(E)=\widetilde{\gamma}_{+}(Q_{1}\cup\cdots\cup Q_{N+1})$ $\lesssim\widetilde{\gamma}_{+}(Q_{1}\cup\cdots\cup Q_{N})+\widetilde{\gamma}_{+}(Q_{N+1})$
$\lesssim\widetilde{\gamma}_{+}(\mathring{Q_{1}}\cup\cdots\cup\mathring{Q}_{N})+\widetilde{\gamma}_{+}(\mathring{Q}_{N+1})$
$\leq 2\widetilde{\gamma}_{+}(\mathring{Q_{1}}\cup\cdots\cup\mathring{Q}_{N+1})=2\widetilde{\gamma}_{+}(\mathring{E}).$
From the previous theorem we can finally deduce the main result of this section:
@@
For any compact set $E\subset\mathbb{R}^{n+1}$,
@@
We may assume $\text{{{A}}}_{\boldsymbol{2}}$ to hold by the outer regularity of $\gamma_{+}$, that by Theorem 3.5 is a comparable capacity to $\widetilde{\gamma}_{+}$. Then, by a direct application of Lemma 6.5 we are done. ∎
@@
The capacity $\widetilde{\gamma}$ is semi-additive. That is, there is an absolute constant $C>0$ so that for any $E_{1},E_{2},\ldots$ disjoint compact sets of $\mathbb{R}^{n+1}$,
@@
In this section we compute the $\widetilde{\gamma}$ capacity of a closed rectangle $R\subset\mathbb{R}^{2}$ with sides parallel to the coordinate axes and respective side lengths $\ell_{x}>0,\,\ell_{t}>0$. More precisely:
@@
We introduce the parameter $r:=\frac{\ell_{x}}{\ell_{t}}$, that is nothing but the spatial side length of $R_{0}$, as well as the measure $\mu:=\pazocal{L}^{2}|_{R_{0}}$.
@@
it follows that for any fixed $t$, $P\ast\mu(x,t)=P\ast\mu(r-x,t),$ or in other words, the one variable function $x\mapsto P\ast\mu(x,t)$ is symmetric with respect to the point $x=r/2$, for any $t\in\mathbb{R}$. Moreover, it is clear that it is nonnegative, tends to 0 as $x\to\pm\infty$ and in fact, for $t>0$, it attains its maximum precisely at $x=r/2$. This last property can be argued as follows: begin by fixing $t\in(0,1]$ and noticing that
@@
$P\ast\mu\bigg{(}\frac{r}{2},t\bigg{)}=\begin{dcases}\text{if }t\leq 0,\hskip 31.2982pt0,\\ \text{if }t\in(0,1],\hskip 14.22636pt\frac{r}{2}\ln\bigg{(}1+\frac{4t^{2}}{r^{2}}\bigg{)}+2t\,\arctan\bigg{(}\frac{r}{2t}\bigg{)},\\ \text{if }t>1,\hskip 31.2982pt\frac{r}{2}\ln\bigg{[}\frac{r^{2}+4t^{2}}{r^{2}+4(t-1)^{2}}\bigg{]}+\,2\arctan\bigg{(}\frac{r}{2(t-1)}\bigg{)}\\ \hskip 113.81102pt-\,2t\bigg{[}\frac{\pi}{2}-\arctan\bigg{(}\frac{2(t-1)}{r}\bigg{)}-\arctan\bigg{(}\frac{r}{2t}\bigg{)}\Bigg{]}.\end{dcases}$
@@
The particular value of $M(r):=P\ast\mu(r/2,1)$ provides a lower bound for $\gamma_{+}$:
@@
Therefore, regarding the original rectangle $R$ and applying Theorem 3.5 and Lemma 6.4, Theorem 7.1 follows. ∎